The measure of an angle is 20 degrees less than three times its supplement. Find the measure of each angle.
I need to find out how to do this proble, my teacher already gave me the answer.

Answer by checkley77(12844)

3X-20+X=180
4X=180+20
4X=200
X=200/4
X=50 DEGREES FOR THE SMALLER ANGLE.
3*50-20=150-20=130 DEGREES FOR THE LARGER ANGLE.
PROOF:
50+130=180
180=180